A Discussion of Oral Media: Urban Legends

Matthew Gray

February 19, 1997

Introduction

Typically, when “media” is discussed, the focus is on technological media, such as the radio, television, films, and recently, software and the World Wide Web. Additionally, “media” is often equated with “mass media” because less well distributed media are less able to have substantial cultural impact. Television, for example, is especially well equipped to very widely deliver a message that has a substantial cultural impact.

Before the existence of these technological innovations, the number of media to choose from was far more restricted. Printed media and oral media still existed, however. In modern culture, it has been argued, because of the economic underpinnings of mass media, cultural narratives have become almost entirely corporate property. While clearly a large amount of modern cultural narrative is generated and owned by large corporations, particularly in the realm of television and films, modern culture is not without narratives that are communally owned. These narratives are primarily transmitted via oral media.

1 A Definition of Oral Media

Oral media here is taken to include all communication that is typically done individual-to-individual, or individual-to-small-group. This is a very inclusive definition, and includes phone calls, email, postal mail, newsletters, community announcements, and any other communication where the people involved know each other personally at some level. This definition is not meant to preclude oral media being a “mass media”, it is merely not centralized mass media.

Urban Legends

The aforementioned narratives are commonly referred to as “Urban Legends”. Urban Legends are typically transmitted via oral media.

Urban Legends typically play to some strong emotion and often, but not always, have some sort of moral message. Usually, for successful transmission, they appeal to some sense of “public good”.

Craig, Mrs. Fields, and the Ankle Slashers

Briefly, presented here are capsule descriptions of three fairly widespread Urban Legends which will be discussed in more detail.

Craig Shergold. Craig Shergold is a poor little British boy dying of cancer and through the “Make-A-Wish” foundation has wished to get in the Guiness Book of World Records for receiving the most Get Well cards. Please send him cards and make a little boy happy.^{1}

Mrs. Fields. A friend of a friend of mine was buying a bunch of Mrs. Fields cookies, and asked the manager if it would be possible to get the recipe. He said no. My friend then asked if she could buy the recipe and was told “Sure, for two fifty”. She had him add it to her bill and took her cookies and recipe home. Upon receiving her credit card bill, she found it was for over $260 dollars, as she had been charged $250 for the recipe. In retaliation for being duped, she is distributing the recipe everywhere, and here it is....^{2}

Ankle Slashers. In mall parking lots across America, a new type of crime is occuring. Young hoodlums will hide underneath cars in parking lots, and when the shoppers return, the hoodlums will slash their ankles from beneath the car, and steal their purchases and sometimes their car.^{3}

Impact of Oral Media

By the fact that these are transmitted orally, they are subject to error and variation more than narratives that are centrally distributed. Variations on each of the above stories exist in substantial variety. For the Craig Shergold story, the boy’s age changes, the boy’s name changes, and whether he wants get well, post, or business cards changes. Variations on the Mrs. Fields story include changes in the recipe, and claiming it is from Niemann Marcus instead of Mrs. Fields. The ankle slashers story is similarly embellished and modified in terms of locale, motivation, and effect.

Additionally, because of the nature of their transmission, these narratives, all of which typically are reported as true, are very difficult to correct when errors are realized. In the latter two cases, these errors are relatively harmless and merely add color to the story, and at worst, cause misplaced emotions toward entities in the stories. In the case of Craig Shergold however, the inability to correct errors has caused substantial problems.

The true story of Craig Shergold is more or less as is related above, but has the added features that the wish was made in 1989, Craig made it into the Guiness Book in 1990, he received so much attention through his wish, an American philanthropist payed for him to recieve medical treatment in the US, and was cured of his cancer in 1991. The mail, however, keeps coming, placing substantial burden on the local Post Offices and the Make-a-Wish Foundation. It is estimated Craig has been sent in excess of 100 million cards, though official counting ceased at 60 million.

Often, corrections succeed in rapidly following the legends, lest someone beleive them true and take rash or inappropriate actions, but the nature of the media makes it very difficult to do so under some circumstances.

Cultural Motivations

Each of these legends has its own reasons for propagation. Spread of the Craig Shergold story seems substantially spurred by a combination of compassion and a sense of being able to relatively easily help in what in many ways seems like a helpless situation. As evidenced by the extent to which this legend has spread, it clearly strikes a particular chord within society: A desire to help, especially when it is easy.

1Entirely true, but incomplete. See Impact of Oral Media section for details

2Entirely false

3Entirely false

The Mrs. Fields story, on the other hand, appeals to a sense of justice. Specifically, it represents a sense of justice that succeeds in vindicating the individual in the face of corporate giants. Once this justice has been served, the story propogates by the desire to share this triumph, and in this case booty, in the form of the recipe.

The ankle slashers story is somewhat less widespread than the others, but is typical of a third brand of fear-motivated legends. Other often recognized legends in this same category include the “laptop thieves at airport security” story, “involuntary organ donation” story, and the “death by jeans shrinking in rain” story. The motivation for distribution of these stories tends to be in the name of enhanced public awareness. In fact most of these stories are entirely false, but the perception of needing to be warned of these threats to our person, particularly in gruesome or otherwise extreme ways, encourages propagation and highlights the specific nature of some the fears society has.

Technological Impact

While each of these legends are relatively modern, and hence seemingly less subject to technological impact, they have in fact come at a time where computer mediated oral communication is becoming widespread.

The Craig Shergold legend, because its basis in fact is well documented, is a case that can be tracked well. The story began its disemmination in 1989, and the actual request was no longer being sent in 1990, and Craig was cured in 1991, all before the recent explosive growth of computer mediated communication. As people began to populate the Internet, substantially driven by the web, new types of oral media became available for distribution of these stories.

The timeline of these events identifies two important points. First, the legend of Craig Shergold spread quite well between 1989 and 1991 via oral media not dominated by email or other computer mediated communication. In fact, in this time over 10 million cards were received.

Second, the widespread use of email spread the story even further. The Make-A-Wish foundation has an 800 number to handle questions about Craig Shergold and noticed a roughly four-fold increase in call volume in 1995, as the Internet craze had recently reached true popular acceptance.^{4}

Conclusion

Culturally owned narratives are still rather ubiquitous even in our mass media society. These narratives are often, however, not recognized as such, but would not likely otherwise be spread. Computer mediated communication is bringing new oral media into popular use, in email, faxes, chat groups, and other discussion forums. Other oral media, however, are still widely and effectively used in society to distribute culturally owned narratives.

4Additionally, there are web sites which fully document the details of these legends, and attempt to dissuade their dissemination, when harmful, as in the case of Craig Shergold